OpenStudy (logan13):

If the diameter of the circle above is 14 mm, what is the area of the circle? 196 mm2 7 mm2 49 mm2 14 mm2

OpenStudy (alexandervonhumboldt2):

radius=1/2 diameter are=radius^2*3.14

OpenStudy (anonymous):

Area = pi * r^2 radius is half of the diameter

OpenStudy (anonymous):

Area = pi * 7^2 Area = 49pi 49 mm^2